Assume the price of good Y rises from Birr 120 to Birr 160 per Kg. and as a result, the consumer demand for good X increases from 50 Kg to 90 Kg. Calculate the cross elasticity of demand of X for Y. 9. Suppose price of a good falls from Birr 12 to Birr 10 per unit, as a result its quantity demanded increases from 100 to 120 units. A) Find price elasticity of demand, and demand for a good is elastic, inelastic, or unit elastic. B) Calculate TR (Total Revenue) at price Birr 12 & 10. C) What is the relationship between TR and price elasticity of demand?


1
Expert's answer
2022-01-13T09:01:35-0500

8. The cross elasticity of demand of X for Y is:

"Ec = \frac{160-120} {90-50} ×\frac{90+50} {160+120} = 0.5."

So, these goods are substitutes.

9.

A) Price elasticity of demand is:

"Ed = \frac{120-100} {10-12} ×\frac{10+12} {120+100} = -1."

So, the demand for a good is unit elastic.

B) TR (Total Revenue) at price Birr 12 & 10 is the same.

C) TR is constant, if price elasticity of demand is -1 (unit elastic).